Plane microwaves are incident on a long slit having a width of 5.0 cm. Calculate the wavelength of the microwaves if the first diffraction minimum is formed at θ = 30°.

Grade:11

1 Answers

Kevin Nash
9 years ago
Sol. For strong transmission, 2 μd = nλ ⇒ λ = 2μd/n Given that, μ = 1.33, d = 1 * 10^-4 cm = 1 * 10^-6 m. ⇒ λ = 2 * 1.33 * 1 * 10^-6/n = 2660 * 10^-9/n m When, n = 4, λ base 1 = 665 nm n = 5, λ base 2 =532 nm n = 6, λ base 3 = 443 nm